OpenStudy (sam_pi):

why the answer to (w^-1/3)/(w^1/4) = 1/w^7/12 ? when I tried to solve it, I subtracted the exponents and I got 1/w^1/12

OpenStudy (photon336):

let's re-write this

OpenStudy (photon336):

\[w^{-\frac{ 1 }{ 3 }} = \frac{ 1 }{ \sqrt[3]{w} }\] do you agree?

OpenStudy (photon336):

@sam_pi

OpenStudy (photon336):

\[ => \frac{ \frac{ 1 }{ \sqrt[3]{w} } }{ \sqrt[4]{w} }\]

OpenStudy (photon336):

\[\frac{ 1 }{ w^{\frac{ 1 }{ 3 }}*w^{\frac{ 1 }{ 4 }} } = \frac{ 1 }{ w^{\frac{ 1 }{ 3 }+\frac{ 1 }{ 4 }} }\]

OpenStudy (photon336):

\[w^{\frac{ 1 }{ 3 }+\frac{ 1 }{4 } = w^{\frac{ 4+3 }{ 12 }} => w^{\frac{ 7 }{ 12 }}}\] \[\frac{ 1 }{ w^{\frac{ 7 }{ 12 }} }\]

OpenStudy (sam_pi):

thanks, I get it now.

OpenStudy (mww):

If you subtract like this: -1/3 - 1/4 = (-4-3)/12 = -7/12 You may have done 1/3 - 1/4 instead which gives 1/12